# Lessons Learned from a Unifying Empirical Study of Parameter-Efficient Transfer Learning (PETL) in Visual Recognition

### v1

@@ A.1 Experiment Details
Throughout all experiments, we employ AdamW optimizer [60] with a batch size of 64 and utilize the cosine decay learning rate scheduler. We train all methods with 100 epochs for VTAB-1K and 40 epochs for many-shot experiments. For the low-shot experiments, the learning rate is tuned from [1e-3, 1e-2] and weight decay from [1e-4, 1e-3]. The method-specific hyperparameter tuning grip is shown in Table 3 with the tunable parameter range in millions. For the many-shot experiment, the learning rate is tuned from [5e-4, 1e-3] and weight decay keeps the same range of [1e-4, 1e-3]. For the robustness experiment, following [92], we set a small learning rate as 3e-5 and weight decay as 5e-3. We use a strong data augmentation following [102].
@@ A.2 Dataset Details
Table 3: Methods-specific hyperparameter tuning grip for VTAB-1K experiment.
@@ B.1 Vision Transformer
Overview of ViT. Inspired by the recent success of Transformer-based models [87] in NLP [91], Vision Transformer (ViT) [17] has become widely used in computer vision. To handle 2D images, ViT divides an image $\bm{I}\in\mathbb{R}^{H\times W\times C}$ into $N$ non-overlapping patches $\{\bm{I}^{(n)}\in\mathbb{R}^{P^{2}\times C}\}_{n=1}^{N}$, where $(H,W)$ is the resolution of the input image, $C$ is the number of channels, $N=HW/P^{2}$ and $(P,P)$ is the resolution of each patch. Each patch $\bm{I}^{(n)}$ is flattened and embedded into a $D$-dimensional vector ${\bm{x}}^{(n)}_{0}$ with a trainable linear projection. Incorporating the BERT design approach [43], a “Class” token ${\bm{x}}_{0}^{(\text{Class})}$ is prepended to the sequence of embedded patches, whose output state at the last Transformer layer is utilized as the image representation. Finally, position embeddings $\mathbf{E}_{\text{pos }}\in\mathbb{R}^{{D\times(1+N)}}$ are added to preserve positional information and form the input $\mathbf{Z}_{0}\in\mathbb{R}^{D\times(1+N)}$ to the ViT, which can be formulated by:
@@
Details of each Transformer layer. As shown in the right part of Figure 8, each Transformer layer consists of a Multi-head Self-Attention (MSA) block, a Multi-Layer Perceptron (MLP) block, and two Layer Normalization (LN) layers [3]. Formally, a Transformer layer $L_{m}$ can be defined as
@@
where $\bm{W}_{1}\in\mathbb{R}^{D\times 4D}$3334 is the MLP ratio in ViT-B, $\bm{W}_{2}\in\mathbb{R}^{4D\times D}$, $b_{1}\in\mathbb{R}^{4D}$, $b_{2}\in\mathbb{R}^{D}$ are weights and biases for two FC layers ($FC_{1}$ and $FC_{2}$) respectively.
@@ B.2 Evaluated Methods
Figure 8: ViT h10, h1, h6 are confusing
@@
B.3 Prompt-based Methods
Prompt-based learning emerged in NLP as an effective approach to adapt pre-trained models for downstream tasks [55, 52]. The core concept involves augmenting the model input with task-specific hints (prompts), which aid the pre-trained model in addressing novel tasks with its existing knowledge. Hard prompts are human-interpretable natural language hints, encompassing task instructions, in-context examples, or supporting information. Alternatively, soft prompts are continuous vector hints that are incorporated into the input embeddings of the input layers or hidden states of other layers. Soft prompts are updated during the fine-tuning process using gradient-based methods, guided by the downstream task-specific loss functions, while the pre-trained model itself remains fixed. The splendent success of prompts in NLP has sparked a growing interest in adopting it in computer vision [97, 86] and multi-modal domains [24].
In this paper, we investigate a prominent and strong prompt-based method called Visual Prompt Tuning (VPT) [39], which represents one of the early endeavours in introducing prompts to computer vision. Specifically, VPT-Shallow adds $l$ prompts $\bm{P}_{0}\in\mathbb{R}^{l\times D}$ to the input of the first Transformer layer $\bm{Z}_{0}$ and the output $\tilde{\bm{P}_{0}}$ of $\bm{P}_{0}$ serves as the input for the next layer as depicted in Equation 11. VPT-Shallow can be perceived as the addition of learnable pixels to the original images. On the other hand, VPT-Deep inserts $l$ prompts $\{\bm{P}_{m}\in\mathbb{R}^{l\times D}\}_{m=0}^{M}$ to the input of every Transformer layer $\bm{Z}_{m}$ but their outputs are discarded at the end of the layer as illustrated in Equation 12.
@@ B.3 Prompt-based Methods
B.4 Adapter-based Methods
Adapter-based methods typically introduce additional trainable parameters into a frozen pre-trained model to facilitate learning of downstream tasks [52]. Initially developed for multi-domain adaptation [74, 75] and continual learning [78, 64], the idea of Adapters is subsequently embraced by Houlsby et al[33] in the NLP domain to adapt Transformer-based networks for downstream tasks, and it also has garnered increasing interest in the computer vision field [97]. In this comparative analysis, we concentrate on five popular Adapter-based methods, encompassing the original Adapter, along with variants focusing on adjusting the positions of Adapters [9, 72], introducing visual inductive biases [40], as well as employing re-parameterization to reduce the number of trainable parameters and inference latency [61].
@@ B.4 Adapter-based Methods
[33] inserts two lightweight bottleneck-structured modules into each Transformer layer: one after the MSA block and the other after the MLP block. As depicted in 9(a), the Adapter is composed of a down-projection layer with $\bm{W}_{\text{down }}\in\mathbb{R}^{r\times D}$, a nonlinear activation function $\sigma$, an up-projection layer with $\bm{W}_{\text{up}}\in\mathbb{R}^{D\times r}$, a scaling factor $s$ and a skip-connection. To limit the number of trainable parameters, the bottleneck dimension is much smaller than the feature dimension $r\ll D$. Formally, Houl. Adapter can be defined as:
@@
[72] is a more efficient variant that introduces the Adapter solely after the MLP block, a strategy that has demonstrated effectiveness in recent studies [34]. Pfeif. Adapter can be defined formally as $h_{9}=\operatorname{Adapter}(h_{9})$ where $\operatorname{Adapter}$ follows Equation 14.
@@
[9] proposed to insert the Adapter in parallel with the MLP block, which differs from the sequential design of Houl. and Pfeif. Adapter. The rationale behind this parallel design lies in the belief that the domain-specific features generated by the Adapter can complement the domain-agnostic features derived from the original MLP block, leading to an improved feature ensemble [83]. Formally, AdaptFormer can be defined as $\bm{h_{9}}=\bm{h_{9}}+\operatorname{Adapter}\bm{(h7)}$ where $\operatorname{Adapter}$ follows Equation 14.
@@
(Convolutional By-Passes) [40] addresses the concern that many existing Adapters lack visual inductive bias, potentially limiting their performance for downstream vision tasks with limited data. To this end, the authors introduce a convolutional bottleneck module, running in parallel with the MSA or(and) MLP block. This module encompasses a $1\times 1$ convolution reducing the channel with $\bm{W}_{\text{down }}\in\mathbb{R}^{r\times D}$, a $3\times 3$ convolution with the same input and output channel, a $1\times 1$ convolution expanding the channel $\bm{W}_{\text{up}}\in\mathbb{R}^{D\times r}$, two nonlinear functions $\sigma$ and a scaling factor $s$, as shown in 9(b). The authors argue that Convpass is more efficient at capturing visual information in low-data scenarios due to its hard-coded locality of convolutional layers. The formal definition of Convpass is shown in Equation 15.
@@
[61] found that the removal of the nonlinear function in the Adapter does not result in performance degradation for vision tasks. In light of this finding, the authors propose a linear Adapter with group-wise transformation [62] and sequentially added two of these linear Adapters to both MSA and MLP blocks. Owing to the sequential placement of the RepAdapter and its inherent linearity, the additional parameters can be re-parameterized to the original MSA or MLP block after training, thereby incurring zero additional costs during inference. RepAdapter is illustrated in 9(c) and formally defined in Equation 16.
@@
where $\bm{W}_{\text{down }}\in\mathbb{R}^{r\times D}$, $\tilde{h}_{g(1,\ldots,G)}\in\mathbb{R}^{\frac{r}{G}\times(N+1)}$ is the features splitted from $\tilde{h}\in\mathbb{R}^{r\times(N+1)}$ and $G$ is the number of groups in group-wise transformation [62]. $\bm{W}_{g(1,\ldots,G)}\in\mathbb{R}^{\frac{D}{G}\times\frac{r}{G}}$ is the projection weight matrix.
@@
B.5 Selective Parameter Tuning Methods
The methods falling within this category aim to selectively update the parameters of a pre-trained model for downstream tasks. Within transfer learning, two prominent strategies, namely full fine-tuning and linear probing [45, 105], represent the two extremes of this category. Full fine-tuning updates all the model parameters end-to-end based on the new dataset while linear probing treats the pre-trained model as a feature extractor and only updates the prediction heads while keeping the backbone frozen. Although full fine-tuning generally exhibits superior performance compared to linear probing [99], it possesses certain limitations that may hinder its practicality in real-world production settings. Firstly, it requires running gradient descent for all parameters and necessitates storing a separate fine-tuned model for each task, incurring significant computational, memory, and storage overhead. These challenges become more salient with Transformer-based models whose parameters grow exponentially. Secondly, full fine-tuning may distort pre-trained features and underperform linear probing in out-of-distribution (OOD) scenarios [47].
To cope with the above issues, a cohort of PETL methods has emerged under this category. In addition to the two common approaches mentioned above, our investigation encompasses seven methods that can be further categorized into two groups: direct selective tuning [98, 5, 93], which involves the direct modification of selective weights, and efficient selective tuning [34, 41, 53], which approximates the weight updates with low-rank factors.
Notably, an extra advantage of methods in this category is that they introduce no additional inference latency, making them particularly favourable when inference efficiency is a priority. Methods within the direct selective tuning group abstain from introducing any new modules, thus inherently avoiding extra inference latency. Meanwhile, for methods in the efficient selective tuning group, the added modules can often be seamlessly integrated into weights of the pre-trained models through the re-parameterization techniques [37, 16], thereby ensuring the absence of increased inference latency as well.
B.5.1 Direct Selective Tuning

### v2

@@ 8 Conclusion
9 Reproducibility Statement
We have made extensive efforts to ensure the reproducibility of our results. All 14 PETL algorithms and two baseline models, along with data preprocessing routines and data loaders for all datasets—including 19 low-shot datasets, 3 many-shot (full) datasets, and 5 robustness-related datasets—are implemented within a systematic and extensible framework. This framework is designed to facilitate the easy addition of new PETL methods and datasets, modification of backbones, and incorporation of additional scenarios, serving as a convenient tool for future research. Detailed explanations of our implementations, raw results for all experiments, and commands to reproduce the results are thoroughly documented in the README file. We provide the anonymous source code in the supplementary material.
10 Ethics Statement
@@ A.1 Experiment Details
VTAB-1K
We employ AdamW optimizer Loshchilov & Hutter (2018) with a batch size of 64 and utilize the cosine decay learning rate scheduler. We train all methods with 100 epochs. The learning rate is tuned from [1e-3, 1e-2] and weight decay from [1e-4, 1e-3]. The method-specific hyperparameter searching grip is shown in Table 3, along with the tunable parameter ranges (in millions). Since most method-specific hyperparameters affect the number of tunable parameters in the PETL methods, we set a cap on the tunable parameters for each PETL method to be less than or equal to 1.5% of the total parameters in ViT-B/16, which is approximately equal to the number of parameters in the Query, Key, and Value matrices of a single MSA block. Consistent with the original VTAB-1k paper (Zhai et al., 2019), most PETL studies Jie & Deng (2023; 2022); Luo et al. (2023); Jia et al. (2022); Zhang et al. (2022b); Luo et al. (2023); Lian et al. (2022) don’t apply data augmentation as it’s challenging to identify a set of augmentations that uniformly benefits all 19 datasets [^4] To demonstrate it, we apply simple data augmentations (RandomResizedCrop and RandomHorizontalFlip) on three datasets in each group, as shown in Table 4. . To ensure that our results are directly comparable and that any performance differences are attributable to the methods themselves rather than data augmentation, we don’t apply data augmentation.
Many-shot
We also employ AdamW optimizer with a batch size of 64 and a cosine decay learning rate scheduler. The learning rate is tuned from [5e-4, 1e-3] and weight decay keeps the same range of [1e-4, 1e-3]. We apply horizontal flipping for CIFAR100, horizontal and vertical flipping for Resisc, and no augmentation for Clevr. We train all methods with 40 epochs.
Robustness Model
CLIP models are trained on image-caption pairs collected from the web. Given a dataset of such pairs $\{(x_{1},s_{1}),\ldots,(x_{B},s_{B})\}$, these models learn an image encoder $g$ and a text encoder $h$ that aim to maximize the similarity $\langle g(x_{i}),h(s_{i})\rangle$ between matching image and caption embeddings while minimizing it for non-matching pairs. For zero-shot classification, the models predict the class of an input image $x$ from a set of $k$ class names $C=\{c_{1},\ldots,c_{k}\}$ by matching $x$ with captions derived from these class names. Specifically, for each class $c_{j}$, a caption is formulated as $s_{j}=\text{``a photo of a }c_{j}\text{''}$ ”. The predicted class is then determined by selecting the one whose caption embedding has the highest similarity with the image embedding: $\hat{y}=\arg\max_{j}\langle g(x),h(s_{j})\rangle$. Alternatively, a weight matrix $W_{\text{zero-shot}}\in\mathbb{R}^{d\times k}$ can be constructed, where each column is the embedding $h(s_{j})$ corresponding to class $c_{j}$. The model’s output scores for each class are then computed as $f(x)=g(x)^{\top}W_{\text{zero-shot}}$. To generate $W_{\text{zero-shot}}$, we ensemble the 80 prompts provided by CLIP at https://github.com/openai/CLIP.
Robustness Setup
In section 3 and section 5, the fine-tuning train set and test set are from the same data distribution. In section 7, we fine-tune the CLIP model using ImageNet-1K training data (100 shots) and subsequently evaluate the fine-tuned model not only on the test set of ImageNet-1K but also on four additional datasets with distribution shifts: ImageNet-V2, ImageNet-R, ImageNet-S, and ImageNet-A, as shown in Figure 8. Following Wortsman et al. (2022), we set a small learning rate as 3e-5 and weight decay as 5e-3. We use a strong data augmentation following Zhang et al. (2022b).
(a) ImageNet Deng et al. (2009)
(b) ImageNetV2 Recht et al. (2019)
(c) ImageNet-R Hendrycks et al. (2021a)
(d) ImageNet Sketch Wang et al. (2019)
(e) ImageNet-A Hendrycks et al. (2021b)
Figure 8: Samples of the class lemon, from the fine-tuned dataset ImageNet and distribution shifts datasets (ImageNet-V2, ImageNet-R, ImageNet-S, and ImageNet-A). The CLIP model is fine-tuned with PETL on ImageNet and evaluated on distribution shifts datasets to measure the robustness of fine-tuned models. The figures are modified based on Wortsman et al. (2022).
Computation
@@ A.2 Dataset Details
Table 3: Methods-specific hyperparameter searching grip for VTAB-1K experiment.
Linear Full VPT-Shallow VPT-Deep BitFit DiffFit LayerNorm SSF
Pfeif.
Adapter
Houl.
Adapter
Adapt-
Former
Rep-
Adapter
Convpass LoRA FacT_TT Fact_TK
Simple DA 84.4 76.8 84.9 84.8 83.8 85.7 85.8 86.1 87.4 86.0 84.9 86.4 85.2 86.8 85.5 86.0
Default 86.6 89.9 88.7 91.5 90.5 90.2 89.7 89.8 91.5 92.1 91.8 92.5 92.1 92.6 91.8 92.5
Caltech101 $\Delta$2.2 -13.1 -3.8 -6.7 -6.7 -4.5 -3.9 -3.7 -4.1 -6.1 -6.9 -6.1 -6.9 -5.8 -6.3 -6.5
Simple DA 67.5 57.8 69.0 71.1 70.7 73.7 73.5 68.4 72.7 72.6 70.6 71.5 71.8 73.0 72.2 71.9
Default 65.7 61.9 67.9 69.4 70.3 71.2 72.2 68.8 72.1 72.3 70.5 69.1 72.0 69.8 71.5 71.8
DTD $\Delta$ 1.8 -4.1 1.1 1.7 0.4 2.5 1.3 -0.4 0.6 0.3 0.1 2.4 -0.2 3.2 0.7 0.1
Simple DA 98.1 92.2 98.2 98.6 98.0 98.8 98.8 98.8 98.4 97.5 98.7 98.0 98.9 98.7 98.7 98.7
Default 98.9 97.4 99.1 99.1 98.9 99.2 99.1 99.1 99.2 98.0 99.2 99.1 99.3 99.1 99.3 99.1
Natural Flower102 $\Delta$0.8 -5.2 -0.9 -0.5 -0.9 -0.4 -0.3 -0.3 -0.8 -0.5 -0.5 -1.1 -0.4 -0.4 -0.6 -0.4
Simple DA 87.3 91.0 88.4 92.0 92.0 91.7 91.9 92.5 92.1 92.5 92.3 93.1 92.7 92.8 93.3 92.8
Default 90.0 88.1 90.3 94.9 95.0 94.1 93.8 94.5 95.5 95.3 95.0 95.3 95.8 94.9 94.9 95.5
EuroSAT $\Delta$2.7 2.9 -1.9 -2.9 -3.0 -2.4 -1.9 -2.0 -3.4 -2.8 -2.7 -2.2 -3.1 -2.1 -1.6 -2.7
Simple DA 74.3 75.0 74.4 80.1 81.0 78.5 80.7 80.6 80.6 81.6 82.2 81.5 81.5 82.2 80.7 82.9
Default 74.9 81.6 77.2 84.2 85.3 80.9 83.0 83.2 85.3 86.5 86.5 86.0 85.9 85.9 85.0 86.0
Resisc45 $\Delta$0.6 -6.6 -2.8 -4.1 -4.3 -2.4 -2.3 -2.6 -4.7 -4.9 -4.3 -4.5 -4.4 -3.7 -4.3 -3.1
Simple DA 74.5 73.6 74.7 76.3 76.3 76.7 76.4 76.4 77.3 75.6 77.0 77.1 76.8 76.2 75.3 77.0
Default 74.6 73.6 74.4 73.9 75.5 75.2 75.2 74.8 76.2 75.2 76.3 75.4 75.9 75.7 75.6 75.7
Specialized Retinopathy $\Delta$0.1 0.0 0.3 2.4 0.8 1.5 1.2 1.6 1.1 0.4 0.7 1.7 0.9 0.5 -0.3 1.3
Simple DA 22.6 29.9 24.3 29.6 28.7 29.0 29.0 27.9 28.9 22.9 30.9 31.4 30.5 30.3 32.5 28.4
Default 29.4 46.6 43.1 56.4 53.9 52.8 52.1 52.1 56.6 54.3 53.0 52.1 55.3 47.2 53.1 53.1
dSpr-Ori $\Delta$6.8 -16.7 -18.8 -26.8 -25.2 -23.8 -23.1 -24.2 -27.7 -31.4 -22.1 -20.7 -24.8 -16.9 -20.6 -24.7
Simple DA 49.8 48.7 49.4 52.7 52.6 54.7 52.7 50.6 53.9 53.6 53.2 52.2 51.3 52.9 52.0 53.3
Default 64.6 77.9 66.5 77.9 79.2 81.0 78.1 81.4 80.2 79.6 80.0 80.2 78.1 79.9 79.3 78.9
KITTI $\Delta$14.8 -29.2 -17.1 -25.2 -26.6 -26.3 -25.4 -30.8 -26.3 -26.0 -26.8 -28.0 -26.8 -27.0 -27.3 -25.6
Simple DA 15.0 24.2 12.8 21.2 21.9 23.4 18.8 24.5 25.1 26.4 24.8 26.9 25.7 24.5 23.4 18.2
Default 17.3 31.0 15.2 33.2 30.1 30.7 24.3 31.9 33.8 34.2 33.0 35.7 38.6 33.4 32.8 27.8
Structured sNORB-Azim $\Delta$2.3 -6.8 -2.4 -12.0 -8.2 -7.3 -5.5 -7.4 -8.7 -7.8 -8.2 -8.8 -12.9 -8.9 -9.4 -9.6
Table 4: We apply simple data augmentations (DA) (RandomResizedCrop and RandomHorizontalFlip) on three datasets in each group. Data augmentation generally does not benefit VTAB-1K and thus, most recent PETL papers (Jie & Deng, 2023; 2022; Luo et al., 2023; Jia et al., 2022; Zhang et al., 2022b; Luo et al., 2023; Lian et al., 2022) skip it. To ensure that our results are directly comparable to existing papers, we don’t apply data augmentation.
@@ B.1 Vision Transformer
Overview of ViT. Inspired by the recent success of Transformer-based models Vaswani et al. (2017) in NLP Wolf et al. (2020), Vision Transformer (ViT) Dosovitskiy et al. (2020) has become widely used in computer vision. To handle 2D images, ViT divides an image $\bm{I}\in\mathbb{R}^{H\times W\times C}$ into $N$ non-overlapping patches $\{\bm{I}^{(n)}\in\mathbb{R}^{P^{2}\times C}\}_{n=1}^{N}$, where $(H,W)$ is the resolution of the input image, $C$ is the number of channels, $N=HW/P^{2}$ and $(P,P)$ is the resolution of each patch. Each patch $\bm{I}^{(n)}$ is flattened and embedded into a $D$-dimensional vector ${\bm{x}}^{(n)}_{0}$ with a trainable linear projection. Incorporating the BERT design approach Kenton & Toutanova (2019), a “Class” token ${\bm{x}}_{0}^{(\text{Class})}$ is prepended to the sequence of embedded patches, whose output state at the last Transformer layer is utilized as the image representation. Finally, position embeddings $\mathbf{E}_{\text{pos }}\in\mathbb{R}^{{D\times(1+N)}}$ are added to preserve positional information and form the input $\mathbf{Z}_{0}\in\mathbb{R}^{D\times(1+N)}$ to the ViT, which can be formulated by:
@@
Details of each Transformer layer. As shown in the right part of Figure 9, each Transformer layer consists of a Multi-head Self-Attention (MSA) block, a Multi-Layer Perceptron (MLP) block, and two Layer Normalization (LN) layers Ba et al. (2016). Formally, a Transformer layer $L_{m}$ can be defined as
@@
where $\bm{W}_{1}\in\mathbb{R}^{D\times 4D}$6664 is the MLP ratio in ViT-B, $\bm{W}_{2}\in\mathbb{R}^{4D\times D}$, $b_{1}\in\mathbb{R}^{4D}$, $b_{2}\in\mathbb{R}^{D}$ are weights and biases for two FC layers ($FC_{1}$ and $FC_{2}$) respectively.
@@ B.2 Evaluated Methods
Figure 9: An overview of a Transformer block in ViT. We adopt the notation $\{h_{1},\ldots,h_{10}\}$ to denote the intermediate features within a Transformer block to facilitate a clearer illustration of the PETL methods discussed in subsection B.2.
@@
B.2.1 Prompt-based Methods
Prompt-based learning emerged in NLP as an effective approach to adapt pre-trained models for downstream tasks Liu et al. (2023); Lialin et al. (2023). The core concept involves augmenting the model input with task-specific hints (prompts), which aid the pre-trained model in addressing novel tasks with its existing knowledge. Hard prompts are human-interpretable natural language hints, encompassing task instructions, in-context examples, or supporting information. Alternatively, soft prompts are continuous vector hints that are incorporated into the input embeddings of the input layers or hidden states of other layers. Soft prompts are updated during the fine-tuning process using gradient-based methods, guided by the downstream task-specific loss functions, while the pre-trained model itself remains fixed. The splendent success of prompts in NLP has sparked a growing interest in adopting it in computer vision Yu et al. (2023b); Tu et al. (2023) and multi-modal domains Gu et al. (2023).
In this paper, we investigate a prominent and strong prompt-based method called Visual Prompt Tuning (VPT) Jia et al. (2022), which represents one of the early endeavours in introducing prompts to computer vision. Specifically, VPT-Shallow adds $l$ prompts $\bm{P}_{0}\in\mathbb{R}^{l\times D}$ to the input of the first Transformer layer $\bm{Z}_{0}$ and the output $\tilde{\bm{P}_{0}}$ of $\bm{P}_{0}$ serves as the input for the next layer as depicted in Equation 11. VPT-Shallow can be perceived as the addition of learnable pixels to the original images. On the other hand, VPT-Deep inserts $l$ prompts $\{\bm{P}_{m}\in\mathbb{R}^{l\times D}\}_{m=0}^{M}$ to the input of every Transformer layer $\bm{Z}_{m}$ but their outputs are discarded at the end of the layer as illustrated in Equation 12.
@@ B.2.1 Prompt-based Methods
B.2.2 Adapter-based Methods
Adapter-based methods typically introduce additional trainable parameters into a frozen pre-trained model to facilitate learning of downstream tasks Lialin et al. (2023). Initially developed for multi-domain adaptation Rebuffi et al. (2017; 2018) and continual learning Rosenfeld & Tsotsos (2018); Mai et al. (2022), the idea of Adapters is subsequently embraced by Houlsby et alHoulsby et al. (2019) in the NLP domain to adapt Transformer-based networks for downstream tasks, and it also has garnered increasing interest in the computer vision field Yu et al. (2023b). In this comparative analysis, we concentrate on five popular Adapter-based methods, encompassing the original Adapter, along with variants focusing on adjusting the positions of Adapters Chen et al. (2022b); Pfeiffer et al. (2021), introducing visual inductive biases Jie & Deng (2022), as well as employing re-parameterization to reduce the number of trainable parameters and inference latency Luo et al. (2023).
@@ B.2.2 Adapter-based Methods
Houlsby et al. (2019) inserts two lightweight bottleneck-structured modules into each Transformer layer: one after the MSA block and the other after the MLP block. As depicted in 10(a), the Adapter is composed of a down-projection layer with $\bm{W}_{\text{down }}\in\mathbb{R}^{r\times D}$, a nonlinear activation function $\sigma$, an up-projection layer with $\bm{W}_{\text{up}}\in\mathbb{R}^{D\times r}$, a scaling factor $s$ and a skip-connection. To limit the number of trainable parameters, the bottleneck dimension is much smaller than the feature dimension $r\ll D$. Formally, Houl. Adapter can be defined as:
@@
Pfeiffer et al. (2021) is a more efficient variant that introduces the Adapter solely after the MLP block, a strategy that has demonstrated effectiveness in recent studies Hu et al. (2021). Pfeif. Adapter can be defined formally as $h_{9}=\operatorname{Adapter}(h_{9})$ where $\operatorname{Adapter}$ follows Equation 14.
@@
Chen et al. (2022b) proposed to insert the Adapter in parallel with the MLP block, which differs from the sequential design of Houl. and Pfeif. Adapter. The rationale behind this parallel design lies in the belief that the domain-specific features generated by the Adapter can complement the domain-agnostic features derived from the original MLP block, leading to an improved feature ensemble Szegedy et al. (2015). Formally, AdaptFormer can be defined as $\bm{h_{9}}=\bm{h_{9}}+\operatorname{Adapter}\bm{(h7)}$ where $\operatorname{Adapter}$ follows Equation 14.
@@
(Convolutional By-Passes) Jie & Deng (2022) addresses the concern that many existing Adapters lack visual inductive bias, potentially limiting their performance for downstream vision tasks with limited data. To this end, the authors introduce a convolutional bottleneck module, running in parallel with the MSA or(and) MLP block. This module encompasses a $1\times 1$ convolution reducing the channel with $\bm{W}_{\text{down }}\in\mathbb{R}^{r\times D}$, a $3\times 3$ convolution with the same input and output channel, a $1\times 1$ convolution expanding the channel $\bm{W}_{\text{up}}\in\mathbb{R}^{D\times r}$, two nonlinear functions $\sigma$ and a scaling factor $s$, as shown in 10(b). The authors argue that Convpass is more efficient at capturing visual information in low-data scenarios due to its hard-coded locality of convolutional layers. The formal definition of Convpass is shown in Equation 15.
@@
Luo et al. (2023) found that the removal of the nonlinear function in the Adapter does not result in performance degradation for vision tasks. In light of this finding, the authors propose a linear Adapter with group-wise transformation Luo et al. (2022) and sequentially added two of these linear Adapters to both MSA and MLP blocks. Owing to the sequential placement of the RepAdapter and its inherent linearity, the additional parameters can be re-parameterized to the original MSA or MLP block after training, thereby incurring zero additional costs during inference. RepAdapter is illustrated in 10(c) and formally defined in Equation 16.
@@
where $\bm{W}_{\text{down }}\in\mathbb{R}^{r\times D}$, $\tilde{h}_{g(1,\ldots,G)}\in\mathbb{R}^{\frac{r}{G}\times(N+1)}$ is the features splitted from $\tilde{h}\in\mathbb{R}^{r\times(N+1)}$ and $G$ is the number of groups in group-wise transformation Luo et al. (2022). $\bm{W}_{g(1,\ldots,G)}\in\mathbb{R}^{\frac{D}{G}\times\frac{r}{G}}$ is the projection weight matrix.
@@
B.2.3 Selective Parameter Tuning Methods
The methods falling within this category aim to selectively update the parameters of a pre-trained model for downstream tasks. Within transfer learning, two prominent strategies, namely full fine-tuning and linear probing Kornblith et al. (2019); Zhuang et al. (2020), represent the two extremes of this category. Full fine-tuning updates all the model parameters end-to-end based on the new dataset while linear probing treats the pre-trained model as a feature extractor and only updates the prediction heads while keeping the backbone frozen. Although full fine-tuning generally exhibits superior performance compared to linear probing Zhai et al. (2019), it possesses certain limitations that may hinder its practicality in real-world production settings. Firstly, it requires running gradient descent for all parameters and necessitates storing a separate fine-tuned model for each task, incurring significant computational, memory, and storage overhead. These challenges become more salient with Transformer-based models whose parameters grow exponentially. Secondly, full fine-tuning may distort pre-trained features and underperform linear probing in out-of-distribution (OOD) scenarios Kumar et al. (2021).
To cope with the above issues, a cohort of PETL methods has emerged under this category. In addition to the two common approaches mentioned above, our investigation encompasses seven methods that can be further categorized into two groups: direct selective tuning Zaken et al. (2022); Basu et al. (2023); Xie et al. (2023), which involves the direct modification of selective weights, and efficient selective tuning Hu et al. (2021); Jie & Deng (2023); Lian et al. (2022), which approximates the weight updates with low-rank factors.
Notably, an extra advantage of methods in this category is that they introduce no additional inference latency, making them particularly favourable when inference efficiency is a priority. Methods within the direct selective tuning group abstain from introducing any new modules, thus inherently avoiding extra inference latency. Meanwhile, for methods in the efficient selective tuning group, the added modules can often be seamlessly integrated into weights of the pre-trained models through the re-parameterization techniques Jacob et al. (2018); Ding et al. (2021), thereby ensuring the absence of increased inference latency as well.
Direct Selective Tuning